# Implementation of tunable frequency-dependent stiffness elements via integrated shunted piezoelectric stacks

### v1

@@ Abstract
Piezoelectric transducers applied on or integrated in structures, combined with appropriate circuits have been extensively investigated as a smart approach to the mitigation of resonant vibrations with high relative amplitudes. Using an approach that is traceable to the use of tuned mass absorbers and dampers, widely used in mechanical and civil engineering, resonantly shunted piezoelectric transducers can be configured to target specific eigenmodes of a structure, if appropriately placed and tuned. By re-framing the function of resonantly shunted piezoelectric transducers as frequency dependent variable stiffness elements, it is of great interest to investigate their capabilities to realize frequency dependent topologies (in the sense of the mechanical connectivity) in the frame of lattice metamaterials, where the load path within a lattice can be interrupted at will for specific frequencies by intercalating null-stiffness components realized with resonantly shunted piezoelectrics. Here, we offer the numerical and experimental verification of this idea, using the example of a potential unit cell of such an adaptive lattice metamaterial. Its realization as additively manufactured component points to the feasibility of such metamaterials in real life.
[…]
@@ 1 Introduction
The implementation of linearly shunted piezoelectric elements in smart structures for vibration reduction is often associated with the surface application of transducers. An inductive shunt in series with a piezo transducer results in an electrical circuit resonance, which interacts with the dynamics of the host structure. This set-up is very well suited for plate-like structures. In correspondence with the large vibrational deformation at a resonant frequency $\omega_{0}$ of the host structure with bending stiffness $K$, the coupling between the shunted transducer and the host structure (Fig. 1) is sufficient to significantly affect the dynamic response of the whole structure in proximity of resonance, as reported by numerous sources [1, 2, 3, 4, 5, 6, 7, 8]. However, the ratio of static stiffness between transducer and host structure is usually quite small. For example, in the cases discussed in [6] and exemplified in Fig. 1 a), the addition of the transducer increases the bending stiffness of the plate by about 9%. The stiffness enhancement stems from the parallel arrangement of $k_{p}$ and $K$ shown in Fig. 1 b) can be calculated from the increase in second moment of area of the bending element, as the bending modes typically correspond to the vibration of interest in this kind of application. Here, we assume that the elastic modulus of the piezoelectric material and the host structure are the same, a good approximation for the case of Lead Zirconate Titanate (PZT) and Aluminum ($E_{PZT}\approx E_{Al}=70GPa$). However, the dynamic stiffness $k_{p}$ of the piezoelectric material is well known to depend on the electrical boundary conditions applied to the transducer. In the case of a resonant shunt, it changes with frequency as shown in 1 c) and reported in [1, 6], with an asymptotic stiffness variation around the $LC$-shunt resonance. Away from the mechanical resonant condition, the variation of the piezoelectric stiffness $k_{p}$, as shown in Fig. 1 c), is hardly discernible in the dynamic response of the structure. Therefore, the parallel structure-piezo arrangement only shows its effect on the dynamic response when the electrical and mechanical resonance frequencies coincide. This parallel arrangement of transducer and plate-like host structure offers significant advantages in terms of structural integrity compared to the series arrangement discussed later, as the strain experienced by the brittle piezoelectric material it limited by the typically very stiff host structure.
@@
Figure 1: a) Surface mounted piezoelectric element, as a typical example of integration strategy for vibration control properties b) representation of the host structure with surface mounted piezoelectric element (as described in [6]) as two spring elements $K$ and $k_{p}$ in parallel, with the piezoelectric as variable stiffness element associated with the variable impedance $Z$ of the electric shunt. c) stiffness change (real part) around the electrical resonance $\omega_{LC}$ of a resonantly (Z=L) shunted transducer, adapted from [9, 1].
@@
Figure 3: (a) The unit cell of structure with integrated piezoelectric stack element, inductively shunted to achieve frequency dependent stiffness. (b) Mesh of the unit cell generated in Ansys for the numerical analysis. The green area shows the constrained surface of the structure.
@@
2 Finite element model of the truss structure with variable stiffness element
A cell of a truss made of aluminium is modelled in Ansys 2020R2. The detailed geometry of the truss, shown in Fig. 3 (b), represents the 3D-printed experimental specimen described in Sec. 3. The diagonal strut, which defines the high rigidity for in-plane deformation, has an embedded PZT cell of 5 mm length. Its properties are given in Tab. 1, and are implemented using Ansys’ Piezoelectric and MEMS toolbox, which also allows to define the top and bottom of the piezo element as electrodes by coupling the voltage of all finite element nodes on these surfaces. The capacitance of the piezo part, as retrieved from the model, is 26.0 pF. This is in line with the expected value for a monolithic piezo crystal with these dimensions. The inductive shunt is implemented using an APDL script. A single node is newly defined, and two CIRCU94 elements are defined between this node and the two electrodes of the PZT element. One element represents the shunt, the other a resistor which allows to change the Q factor of the shunt.
Table 1: Properties of the PZT element in the model
[…]
@@ 2 Finite element model of the truss structure with variable stiffness element
$L=\frac{1}{\omega_{0}^{2}C}$ (3)
are thus 26.7 H, 10.4 H and 7.14 H. The Q-factor of the shunt can be regulated by the series resistor through
$Q=\frac{\omega_{0}L}{R},$ (4)
@@
The potential of the piezo stack to act as a tunable stiffness element is shown by detuning the shunt to an inductance value leading to an electrical resonance which does not coincide with a mechanical resonance. This is conventionally not of much interest, since low vibration amplitudes lead to small changes in the overall stiffness, especially in the parallel transducer arrangement. However, given the axial stress in the diagonal strut and the resonant behaviour of the shunt, this regime for the series arrangement has a measurable effect. As illustrated in Fig. 4, a shunt of 17.1 H induces a new mechanical resonance at 7.6 kHz. The piezo element loses its mechanical stiffness at the resonance frequency, thereby compromising the rigidity of the entire diagonal strut. The resulting deformation is visualized in 5 (d): the piezo element loses its dynamic stiffness, thereby allowing large axial deformations of the diagonal strut, and a large overall deformation of the truss structure.
Figure 5: Modal deformation of the truss at 6.0 kHz (a) and 9.7 kHz (b). The first mode is characterized by bending of the diagonal strut, the second mode by a deformation along the axis of the strut. The color scale shows normalized vibration velocities. Deformed shape at 7.6 kHz with identical scaling factor ×$\times 10^{6}$ for the visualization of the deformation, in the open-circuit (c) and shunted (d) case. The overall deformation of the truss, and the extreme compression of the piezo due to a loss of dynamic stiffness are clear.
@@ 3 Experimental validation of the variable stiffness structure
The experimental setup is shown in Fig. 6. The sample is mounted on an electrodynamic shaker by means of a force sensor (PCB208C01), whose output is used as reference signal for the FRF calculations. The response of the truss is measured on the vertical strut opposite the excitation point (at the bottom of the figure) by a scanning laser vibrometer (Polytec PSV-400). For practical reasons, the inductance is realized by an Antoniou synthetic gyrator circuit, powered by 2 batteries. This classical electrical circuit using two operational amplifiers can be used as a tunable inductor, by varying the resistance value of a potentiometer. Apart from being able to generate inductances in the desired range, the $Q$-factor can be kept at a sufficiently high value to achieve a measurable variable stiffness.
Figure 7: Measured frequency response function, averaged over a vertical strut, of the bracket for 4 different inductive shunts. The black dashed line shows the unshunted case. The vertical dashed lines show the electrical shunt resonance frequencies $\omega_{0}=1/\sqrt{LC}$. The shunt tuned to the first eigenmode does not have a measurable effect, whereas correct tuning reduces the modal amplitude of the second and third peak. Mistuning the shunt leads to an additional mechanical resonance.
The experimental results are in line with the simulations, although the resonance frequencies show some deviation. This is due to geometrical differences, and unknown stiffness and damping in the connection to the shaker. For correctly tuned inductance values, the second and third eigenmodes are efficiently suppressed, whereas the first eigenmode is not affected. For an inductance leading to an electrical resonance that does not coincide with a mechanical eigenfrequency, a clear but rather feeble additional mechanical resonance can be noticed. The $Q$-factor of the Antoniou circuit is too low to achieve more pronounced effects. It is remarkable, however, that even far away from mechanical resonances the electromechanical coupling between the structure and the embedded piezo can change the dynamic stiffness.
As is well known, changing the $Q$-factor of the shunt can have additional beneficient effects on the modal damping. This is illustrated in Fig, 8 for the damping of the third eigenmode.
@@ 5 Conclusions and outlook
The experiments and the corresponding numerical models confirm that the increased electromechanical coupling, associated with a series arrangement of the piezo element within a vibrating structure, allows for a modification of the dynamic response that goes beyond the modal vibration reduction demonstrated in [10, 11, 12]. The excellent coupling allows for a full loss of stiffness in the diagonal strut of a truss structure, as demonstrated by the introduction of additional vibration modes stemming from the modification of the stiffness of the diagonal strut. The tunability of the system by using an Antoniou-gyrator inductance gives full control over the vibrational properties of the system. At the level of the individual unit shown in Fig. 3, the loss of stiffness can for instance be beneficial to avoid damage at high loads by letting the structure yield, or to create vibrating elements that can be used as tuned vibration absorbers. From the perspective of the synthesis of novel architected materials that transcend the example discussed in [14], the ability to modify at will the connectivity of a lattice structures is a very appealing way to create lattice structures with frequency dependent unit cell geometry. Here, the factor limiting the implementation of this tunable unit cell geometry scheme over a wide range of frequencies, is the range of stable inductances and fairly low $Q$-factors that can customarily be achieved by Antoniou circuit implementation of synthetic gyrators. Alternative synthetic gyrator designs may offer relief to this constraint.

### v2

@@ Abstract
Piezoelectric transducers applied on or integrated in structures, combined with appropriate circuits have been extensively investigated as a smart approach to the mitigation of resonant vibrations with high relative amplitudes. A resonant shunt circuit consisting of the capacitive piezoelectric transducer and an inductance can be configured to target specific eigenmodes of a structure, if appropriately placed and tuned. Their effect is expressed in terms of mechanical impedance of the host structure, allowing for the exchange of energy between the mechanical and electrical domain, to dramatically affect the dynamic response of the structure. By re-framing the function of resonant shunted piezoelectric transducers as frequency dependent variable stiffness elements, this paper investigates their capability to realize a frequency dependent structural mechanical connectivity, where the load path within a lattice structure can be interrupted at will for specific frequencies by tunable null-stiffness components. Here, we offer the numerical and experimental verification of this idea, by demonstrating the ability to significantly affect the dynamic response of a unit cell of an adaptive lattice metamaterial, even away from a structural resonance. In the latter case, the null-stiffness shunt leads to an additional resonance peak in the truss’ dynamic response. Its realization as additively manufactured component points to the feasibility of such structures in real life.
: Smart Mater. Struct.
@@ 1 Introduction
The implementation of linearly [1, 2, 3, 4] or non-linearly [5] shunted piezoelectric elements in smart structures for vibration reduction is often associated with the surface application [6, 7] of transducers onto plate like structures. An inductive shunt in series with a piezo transducer results in an electrical circuit resonance, which interacts with the dynamics of the host structure. This set-up is very well suited for plate-like structures. In correspondence with the large vibrational deformation at a resonant frequency $\omega_{0}$ of the host structure with bending stiffness $K$, the coupling between the shunted transducer and the host structure (Fig. 1) is sufficient to significantly affect the dynamic response of the whole structure in proximity of resonance, as reported by numerous sources [8, 9, 10, 11, 12, 13, 14, 15]. However, the ratio of static stiffness between transducer and host structure is usually quite small. For example, in the cases discussed in [13] and exemplified in Fig. 1 a), the addition of the transducer increases the bending stiffness of the plate by about 9%. The stiffness enhancement stems from the parallel arrangement of $k_{p}$ and $K$ shown in Fig. 1 b) can be calculated from the increase in second moment of area of the bending element, as the bending modes typically correspond to the vibration of interest in this kind of application. Here, we assume that the elastic modulus of the piezoelectric material and the host structure are the same, a good approximation for the case of Lead Zirconate Titanate (PZT) and Aluminum ($E_{PZT}\approx E_{Al}=70GPa$). However, the dynamic stiffness $k_{p}$ of the piezoelectric material is well known to depend on the electrical boundary conditions applied to the transducer. In the case of a resonant shunt, it changes with frequency as shown in 1 c) and reported in [8, 13], with an asymptotic stiffness variation around the $LC$-shunt resonance. Away from the mechanical resonant condition, the variation of the piezoelectric stiffness $k_{p}$, as shown in Fig. 1 c), is hardly discernible in the dynamic response of the structure. Therefore, the parallel structure-piezo arrangement only shows its effect on the dynamic response when the electrical and mechanical resonance frequencies coincide. This parallel arrangement of transducer and plate-like host structure offers significant advantages in terms of structural integrity compared to the series arrangement discussed later, as the strain experienced by the brittle piezoelectric material it limited by the typically very stiff host structure.
@@
Figure 1: a) Surface mounted piezoelectric element, as a typical example of integration strategy for vibration control properties b) representation of the host structure with surface mounted piezoelectric element (as described in [13]) as two spring elements $K$ and $k_{p}$ in parallel, with the piezoelectric as variable stiffness element associated with the variable impedance $Z$ of the electric shunt. c) stiffness change (real part) around the electrical resonance $\omega_{e}$ of a resonantly (Z=L) shunted transducer, adapted from [16, 8].
@@
Figure 3: (a) The unit cell structure with integrated piezoelectric stack element, inductively shunted to achieve frequency dependent stiffness. (b) Mesh of the unit cell generated in Ansys for the numerical analysis. The green area shows the constrained surface of the structure.
@@
2 Finite element modeling of the variable stiffness element
2.1 Variable stiffness of the piezo stack with shunt
The interaction between an inductively shunted piezo patch and the host structure can be explained in function of a change in mechanical impedance, as described by [8]. The impedance ratio between the shunted and open-circuit piezo patch is given by the formula
$Z=1-k^{2}_{33}\left(\frac{\delta^{2}}{\gamma^{2}+\delta^{2}r\gamma+\delta^{2}}\right),$ (3)
where $\delta=\omega_{e}/\omega_{n}$ is the ratio between the electrical shunt frequency $\omega_{e}=1/\sqrt{LC}$ and the mechanical natural frequency of the piezo patch. The normalized frequency variable is $\gamma=i\omega/\omega_{n}$, and $r=RC\omega_{n}$ describes the damping due to the resistance in the shunt. Finally,
$k_{33}=\frac{d_{33}}{\sqrt{s_{33}\varepsilon_{3}}}$ (4)
includes the electromechanical coupling properties of the piezo patch, with $d_{33}$ being the piezoelectric coupling coefficient (in C/N), $s_{33}$ the mechanical compliance (the inverse of the Young’s modulus), and $\varepsilon_{3}$ the electrostatic permittivity, in the corresponding directions.
For the case described in this paper, where a piezo material is integrated in the mechanical structure to change its overall dynamic stiffness, a high reduction of the mechanical impedance is required. At first, it is verified that a finite element model of the shunted piezoelectric patch behaves as predicted by theory. The model is constructed in Ansys 2020R2. Its properties are given in Tab. 1, and are implemented using Ansys’ Piezoelectric and MEMS toolbox, which also allows to define the top and bottom of the piezo element as electrodes by coupling the voltage of all finite element nodes on these surfaces. The capacitance of the piezo part, as retrieved from the model, is 22.7 pF. This is in line with the expected value for a monolithic piezo crystal with these dimensions. The inductive shunt is implemented using an APDL script. A single node is newly defined, and two CIRCU94 elements are defined between this node and the two electrodes of the PZT element. One element represents the shunt, the other a resistor which allows to change the Q factor of the shunt. For two cases, open circuit and shunted, a harmonic unit force is applied on the top electrode while the bottom side of the piezo is rigidly blocked (Fig. 4 (a)). The complex frequency response function of the displacement of the top electrode is extracted between 6 and 16 kHz, so that the mechanical complex stiffness can be calculated.
Table 1: Properties of the piezoelectric element in the model
Density 7800 kg/m3
@@ 2.1 Variable stiffness of the piezo stack with shunt
The real part and absolute value of the dynamic stiffness are shown in Fig. 4 (b) and (c). An open-circuit simulation shows that the first longitudinal mode of the piezo material occurs at $\omega_{n}=200$ kHz, which is far higher than the desired inductive shunt eigenfrequencies. A characteristic decrease followed by an increase of the mechanical impedance can be noticed, with only a small difference between the analytical and numerical response, as shown in Fig. 3 (b). This difference is due to the orthotropic mechanical and piezoelectric properties included in the finite element model, which are not captured in the analytical model. Fig. 3 (d) illustrates the mechanical stiffness reduction as a function of the shunt’s Q-factor. For values higher than 50, a more than 90% reduction can be seen, whereas the stiffness change decreases rapidly with decreasing Q-values.
2.2 Truss structure with variable stiffness element
A cell of a truss made of aluminium is modelled in Ansys 2020R2. The detailed geometry of the truss, shown in Fig. 3 (b), represents the 3D-printed experimental specimen described in Sec. 3. The diagonal strut, which defines the high rigidity for in-plane deformation, has an embedded PZT cell of 5 mm length.
A harmonic analysis is performed on the cell without an inductive shunt: the top right corner is elastically supported, and the entire structure is excited by a 10 m/s2 acceleration to mimic a body force. This reflects the experimental conditions, where the light structure is entirely excited by an electrodynamic shaker. The truss cell has 3 in-plane vibrational modes in the frequency range from 5 to 13 kHz, at roughly 6, 9.5, and 11.5 kHz. Other vibration modes, i.e. the modes with deformation shapes normal to the truss plane, are not considered for this study since the diagonal strut does not play a crucial role.
@@ 2.2 Truss structure with variable stiffness element
$L=\frac{1}{\omega_{e}^{2}C}$ (5)
are thus 28.0 H, 11.8 H and 8.4 H. The Q-factor of the shunt can be regulated by the series resistor through
$Q=\frac{\omega_{e}L}{R},$ (6)
@@
The potential of the piezo stack to act as a tunable stiffness element is shown by detuning the shunt to an inductance value leading to an electrical resonance which does not coincide with a mechanical resonance. This is conventionally not of much interest, since low vibration amplitudes lead to small changes in the overall stiffness, especially in the parallel transducer arrangement. However, given the axial stress in the diagonal strut and the resonant behaviour of the shunt, this regime for the series arrangement has a measurable effect. As illustrated in Fig. 5, a shunt of 19.8 H induces a new mechanical resonance at 7.6 kHz. The piezo element loses its mechanical stiffness at the resonance frequency, thereby compromising the rigidity of the entire diagonal strut. The resulting deformation is visualized in 6 (d): the piezo element loses its dynamic stiffness, thereby allowing large axial deformations of the diagonal strut, and a large overall deformation of the truss structure.
Figure 6: Modal deformation of the truss at 6.0 kHz (a) and 9.7 kHz (b). The first mode is characterized by bending of the diagonal strut, the second mode by a deformation along the axis of the strut. The color scale shows normalized vibration velocities. Deformed shape at 7.6 kHz with identical scaling factor ×$\times 10^{6}$ for the visualization of the deformation, in the open-circuit (c) and shunted (d) case. The overall deformation of the truss, and the extreme compression of the piezo due to a loss of dynamic stiffness are clearly noticeable.
@@ 3 Experimental validation of the variable stiffness structure
The experimental setup is shown in Fig. 7. The sample is mounted on an electrodynamic shaker by means of a force sensor (PCB208C01), whose output is used as reference signal for the FRF calculations. The response of the truss is measured on the vertical strut opposite the excitation point (at the bottom of the figure) by a scanning laser vibrometer (Polytec PSV-400). The excitation signal is a sine sweep ranging from 4 to 13 kHz. This spans the first three in-plane resonance frequencies. Higher frequencies could not be achieved by the shaker system, and at lower frequencies the combined system of the shaker and sample yields rigid body modes of the sample. The latter are not relevant to the study at hand, since the struts are not deformed and no effect of the piezo stack can be expected. For practical reasons, the inductance is realized by an Antoniou synthetic gyrator circuit [24], powered by 2 batteries. This classical electrical circuit using two operational amplifiers can be used as a tunable inductor, by varying the resistance value of a potentiometer. Apart from being able to generate inductances in the desired range, the $Q$-factor can be tuned by an additional potentiometer in series. It can be kept at a sufficiently high value to achieve a measurable variable stiffness.
Figure 8: Measured frequency response function, averaged over a vertical strut, of the bracket for 4 different inductive shunts. The black dashed line shows the unshunted case. The vertical dashed lines show the electrical shunt resonance frequencies $\omega_{e}=1/\sqrt{LC}$. The shunt tuned to the first eigenmode does not have a measurable effect, whereas correct tuning reduces the modal amplitude of the second and third peak. Mistuning the shunt leads to an additional mechanical resonance.
The experimental results shown in Fig. 8 are in line with the simulations, showing three main resonance peaks in the chosen frequency range, and similar velocity amplitudes. The resonance frequencies show some deviation. This is due to geometrical differences, and the unknown stiffness and damping of the connection to the shaker, leading to complex boundary conditions that can only be approximately reproduced in the model. Moreover, the experimental spectrum is richer than the modeled spectrum showing several minor peaks because of the inevitable out-of-plane vibrations of the structure. For correctly tuned inductance values, the second and third eigenmodes are efficiently suppressed, whereas the first eigenmode is not affected. For an inductance leading to an electrical resonance that does not coincide with a mechanical eigenfrequency, a clear but rather feeble additional mechanical resonance can be noticed. The $Q$-factor of the Antoniou circuit is too low to achieve more pronounced effects. It is remarkable, however, that even far away from mechanical resonances the electromechanical coupling between the structure and the embedded piezo can change the dynamic stiffness.
As is well known, changing the $Q$-factor of the shunt can have additional advantageous effects on the modal damping. This is illustrated in Fig, 9 for the damping of the third eigenmode.
@@ 5 Conclusions and outlook
The experiments and the corresponding numerical models confirm that the increased electromechanical coupling, associated with a series arrangement of the piezo element within a vibrating structure, allows for a modification of the dynamic response that goes beyond the modal vibration reduction demonstrated in [17, 18, 19]. The excellent coupling allows for a full loss of stiffness in the diagonal strut of a truss structure, as demonstrated by the introduction of additional vibration modes stemming from the modification of the stiffness of the diagonal strut. However, the embedded piezoelectric stack is only effective if it undergoes a net strain along its polarized axis. A pure bending deformation, as is seen in the first vibration mode of the truss, does not lead to a measurable effect on the dynamics of the system.
For future applications, the size and load bearing capacity of the piezoelectric stacks will be defining factors. Although larger stacks can undergo larger deformations, have a higher capacitance, and should therefore be able to affect lower frequencies for reasonable inductance values, they are not commercially available in sizes larger than several tens of millimeters. Moreover, to achieve tunable stiffness of structures with a more complex, potentially three-dimensional, shape, the placement of the stack has to be well considered. If the local strain is not sufficiently high, the effect on the structure’s dynamics will be negligible.
The tunability of the system by using an Antoniou-gyrator inductance gives full control over the vibrational properties of the system. At the level of the individual unit shown in Fig. 3, the loss of stiffness can for instance be beneficial to avoid damage at high loads by letting the structure yield, or to create vibrating elements that can be used as tuned vibration absorbers. From the perspective of the synthesis of novel architected materials that transcend the example discussed in [21], the ability to modify at will the connectivity of a lattice structures is a very appealing way to create lattice structures with frequency dependent unit cell geometry. Here, the factor limiting the implementation of this tunable unit cell geometry scheme over a wide range of frequencies, is the range of stable inductances and fairly low $Q$-factors that can customarily be achieved by Antoniou circuit implementation of synthetic gyrators. Alternative synthetic gyrator designs may offer relief to this constraint.